The Tree House has a pretax cost of debt of 6.7 percent and a return on assets of 10.4 percent. The debt–equity ratio is .47. Ignore taxes. What is the cost of equity? Multiple Choice 12.75% 13.24% 12.53% 8.66% 12.14%
Answer #1

debt–equity ratio=debt/equity

Hence debt=0.47 equity

Let equity be $x

Debt=$0.47 x

Total=$1.47 x

Return on assets=Respective costs*Respective weight

10.4=(x/1.47x*Cost of equity)+(0.47x/1.47x*6.7)

10.4=(1/1.47*Cost of equity)+2.142176871

Cost of equity=(10.4-2.142176871)*1.47

=12.14%(Approx).